OpenStudy (anonymous):

use the comparison limit sum_{n=1}^\infty {8 n^{11} - n^{7} + 7\sqrt{n}}/{5 n^{13} - n^{6} + 5 }

OpenStudy (agreene):

\[\sum_{n=1}^\infty \frac{8 n^{11} - n^{7} + 7\sqrt{n}}{5 n^{13} - n^{6} + 5 }\] yes?